Mass and Energy Balances

Mass and Energy Balances – p.1/6

Why, a four-year-old child could understand this. Someone get me a four-year-old child

Groucho Marx

Basic Steps in Flowsheet Synthesis

Introduction

✓Gather information about the process chemistry ✓Generate flow diagram based on Douglas Hierarchy Solve mass and energy balances Estimate equipment size based on flow rates from previous step Estimate equipment cost based on size from previous step Optimize process We developed a sketch of a process in the previous lecture.

The next step is to determine the size of each unit, which will help us estimate the cost of each unit.

There are two ways to do this:

1. The Chemist’s Approach: Build larger sizes of laboratory equipment and experimentally measure all the process variables.

2. The Engineer’s Approach: Develop process models for each unit and solve these equations to estimate the mass and energy flow rates. Use these rates to determine size.

We developed a sketch of a process in the previous lecture.

The next step is to determine the size of each unit, which will help us estimate the cost of each unit.

There are two ways to do this:

1. The Chemist’s Approach: Build larger sizes of laboratory equipment and experimentally measure all the process variables.

2. The Engineer’s Approach: Develop process models for each unit and solve these equations to estimate the mass and energy flow rates. Use these rates to determine size.

We developed a sketch of a process in the previous lecture.

The next step is to determine the size of each unit, which will help us estimate the cost of each unit.

There are two ways to do this:

1. The Chemist’s Approach: Build larger sizes of laboratory equipment and experimentally measure all the process variables.

2. The Engineer’s Approach: Develop process models for each unit and solve these equations to estimate the mass and energy flow rates. Use these rates to determine size.

We developed a sketch of a process in the previous lecture.

The next step is to determine the size of each unit, which will help us estimate the cost of each unit.

There are two ways to do this:

1. The Chemist’s Approach: Build larger sizes of laboratory equipment and experimentally measure all the process variables.

2. The Engineer’s Approach: Develop process models for each unit and solve these equations to estimate the mass and energy flow rates. Use these rates to determine size.

Process Models

A process model is a set of equations, including the necessary input data to solve the equations, that allow us to predict the behavior of a chemical process system.

1. Fundamental Model: Derived from Material and energy balances Physical and Chemical Properties

2. Empirical Model: Obtained by “fitting” data Numerical Methods Statistical Analysis A process model is a set of equations, including the necessary input data to solve the equations, that allow us to predict the behavior of a chemical process system.

1. Fundamental Model: Derived from Material and energy balances Physical and Chemical Properties

2. Empirical Model: Obtained by “fitting” data Numerical Methods Statistical Analysis A process model is a set of equations, including the necessary input data to solve the equations, that allow us to predict the behavior of a chemical process system.

1. Fundamental Model: Derived from Material and energy balances Physical and Chemical Properties

Types of Process Models

2. Empirical Model: Obtained by “fitting” data Numerical Methods Statistical Analysis 1. Steady state models.

2. Dynamic models.

In this course, we will develop:

steady state models from first principles. approximate models of chemical processes.

Dynamic models will be considered in ECH 4323: Chemical Process Control.

1. Steady state models.

2. Dynamic models.

In this course, we will develop:

steady state models from first principles. approximate models of chemical processes.

Dynamic models will be considered in ECH 4323: Chemical Process Control.

1. Steady state models.

2. Dynamic models.

In this course, we will develop:

steady state models from first principles. approximate models of chemical processes.

Dynamic models will be considered in ECH 4323: Chemical Process Control.

In particular, we will consider the following unit operations: Mixer and Splitter Reactor Flash Distillation Absorption Multicomponent Systems Linear Balances

Models for other unit operations may be found in:

Transport Processes and Separation Process Principles by Geankoplis Separation Process Principles by Seader and Henley Elements of Chemical Reaction Engineering by Fogler In particular, we will consider the following unit operations: Mixer and Splitter Reactor Flash Distillation Absorption Multicomponent Systems Linear Balances

Models for other unit operations may be found in:

Transport Processes and Separation Process Principles by Geankoplis Separation Process Principles by Seader and Henley Elements of Chemical Reaction Engineering by Fogler In particular, we will consider the following unit operations: Mixer and Splitter Reactor Flash Distillation Absorption Multicomponent Systems Linear Balances

Models for other unit operations may be found in:

Transport Processes and Separation Process Principles by Geankoplis Separation Process Principles by Seader and Henley Elements of Chemical Reaction Engineering by Fogler In particular, we will consider the following unit operations: Mixer and Splitter Reactor Flash Distillation Absorption Multicomponent Systems Linear Balances

Models for other unit operations may be found in:

Transport Processes and Separation Process Principles by Geankoplis Separation Process Principles by Seader and Henley Elements of Chemical Reaction Engineering by Fogler